### v1

# On orthogonal decompositions of hermitian Higgs bundles

Abstract.
An hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle [15] and $(E,h)$ is a holomorphic hermitian vector bundle [12]. As it is well known, several results on holomorphic vector bundles can be extended to the Higgs bundles setting. In this article we show that some classical propositions, involving orthogonal decompositions of holomorphic hermitian vector bundles and the second fundamental form of its holomorphic subbundles, can be extended to hermitian Higgs bundles. The extended results concerning orthogonal decompositions have immediate applications in Higgs bundles, and we include them in the form of two corollaries. The extended results concerning the second fundamental form are previously known in literature [2]. Here we revisit these results and include alternative proofs without using local computations.
Keywords: Higgs bundle; Kähler manifold; hermitian metric.
2010 Mathematics Subject Classification:
Primary 53C07, 53C55, 32C15; Secondary 14J60, 32G13
1. Introduction
Roughly speaking a Higgs bundle is a holomorphic vector bundle, together with a section of a certain associated holomorphic vector bundle. Since its introduction by Hitchin [9] and Simpson [15, 16], these geometric objects have played an important role in complex geometry as well as in mathematical-physics. For instance, from the viewpoint of geometry, several results on holomorphic vector bundles concerning special metrics and connections can be extended to Higgs bundles [2, 3, 4, 5, 13, 14]. Similarly, from the viewpoint of physics, Higgs bundles are closely related to dimensional reduction. In fact, in the pioneering work [9], Higgs bundles arise as geometric objects associated to Hitchin’s equations (a two dimensional reduction of the self-dual Yang-Mills equations in four dimensions). On the other hand, Hitchin’s equations can also be seen as the lowest dimensional case of another set of equations introduced by Ward [19] and which are a dimensional reduction of Yang-Mills equations in higher dimensions. The equations introduced by Ward are known as $2k$-Hitchin’s equations and can be directly related to Higgs bundles [6]. It is important to mention that from the very beginning, Higgs bundles are associated to another set of equations, usually known as the Hermitian-Yang-Mills equations [15]. Higgs bundles are to HYM equations as holomorphic vector bundles are to Hermitian-Einstein’s equations [12, 17]. Today, Higgs bundles and all the above equations are of major interest in geometry and physics and there exists a very ample literature related to the topic, see for instance [1, 11, 18, 20]. In particular, Higgs bundles are geometric objects of relevance in the foundational work of Kapustin and Witten [10].
This article is organized as follows. In Section 1 we review part of the elementary theory on holomorphic bundles that will be important later on. In particular, we explicitly recall four classical propositions on holomorphic hermitian vector bundles. In this section we also establish part of the notation and terminology that will be used throughout the article. In Section 2 we summarize some basic concepts and properties on Higgs bundles. We also mention two lemmas concerning short exact sequences and invariance conditions on Higgs bundles that will be relevant in the remaining part of the article. In this section we also revisit some propositions on Higgs bundles [2], which are extensions of well known results associated to the second fundamental form of holomorphic subbundles of holomorphic hermitian vector bundles [12]. We include here alternative proofs of these results that do not use local computations. In Section 3 we establish a couple of propositions concerning orthogonal decompositions of hermitian Higgs bundles. In particular, we include proofs of these results that rely heavily on classical results on holomorphic hermitian vector bundles as well as standard definitions on Higgs bundles. Finally, in this section we show some immediate applications on Higgs bundles of the results previously established. More specifically, we show that using the aforementioned lemmas and by introducing the notion of $C^{\infty}$ Higgs orthogonal complement, the extended results on Higgs bundles involving orthogonal decompositions imply two corollaries, which can be established by referring exclusively to objects in the Higgs category.
The main purpose of this article is to establish a couple of propositions as well as to revisit two other results on hermitian Higgs bundles [2]. The first propositions are natural extensions of well known propositions on orthogonal decompositions of holomorphic hermitian vector bundles. These classical results are included in the form of propositions (4.18) and (6.14) in the first chapter of the celebrated textbook of Kobayashi [12]. There are two other classical results, closely related to the previous propositions and which appear in the same reference as propositions (6.4) and (6.6). Extensions to hermitian Higgs bundles of these latter results have been previously studied in [2]. For the reader convenience, in this section we review some elementary definitions as well as summarize all these classical results on holomorphic vector bundles (for more details the reader can see [7, 8, 12]).
Throughout this article, $M$ is an $n$-dimensional compact Kähler manifold and $E$ is a rank $r$ holomorphic vector bundle over it. We denote by $\Omega^{1,0}$ and $\Omega^{0,1}$ the holomorphic cotangent bundle to $M$ and its complex conjugate bundle and by $\Omega^{p,q}$, with $0\leq p,q\leq n$, the $C^{\infty}$ complex vector bundle on $M$ obtained by taking wedge products ($p$ and $q$ times) of $\Omega^{1,0}$ and $\Omega^{0,1}$. If $1\leq s\leq 2n$ we have $\Omega^{s}=\oplus_{p+q=s}\Omega^{p,q}$ and we denote by $A^{s}$ and $A^{p,q}$ (resp. $A^{s}(E)$ and $A^{p,q}(E)$) the spaces of $C^{\infty}$ complex $s$-forms and $(p,q)$-forms on $M$ (resp. forms on $M$ with coefficients in $E$), i.e., they are the spaces of $C^{\infty}$ sections of $\Omega^{s}$ and $\Omega^{p,q}$ (resp. of $\Omega^{s}\otimes E$ and $\Omega^{p,q}\otimes E$). As it is known, any hermitian metric $h$ in $E$ determines a unique connection $D_{h}$ in $E$, usually known as the Chern or the hermitian connection of $h$. Following the notation in [12] we decompose $D_{h}=D^{\prime}_{h}+d^{\prime\prime}$ into its $(1,0)$ and $(0,1)$ parts. The curvature of this connection is given by $R_{h}=D_{h}\wedge D_{h}$ and is an ${\rm End}E$-valued form of type $(1,1)$, usually called the Chern or the hermitian curvature of $h$. Hence, $R_{h}$ is an element in $A^{1,1}({\rm End}E)$. Using the above decomposition in types of $D_{h}$ we have
$R_{h}=D^{\prime}_{h}\wedge d^{\prime\prime}+d^{\prime\prime}\wedge D^{\prime}_{h}\,.$
Given an $h$ in $E$, the pair $(E,h)$ is called a holomorphic hermitian vector bundle and it is common to say that $D_{h}$ and $R_{h}$ are the Chern connection and curvature of $(E,h)$, respectively. Clearly, we can consider $C^{\infty}$ complex subbundles of $E$ and not only holomorphic subbundles. In particular, if $E^{\prime}\subset E$ is a $C^{\infty}$ complex subbundle we define the orthogonal complement $E^{\prime\prime}$ of $E^{\prime}$ with respect to $h$, which is also a $C^{\infty}$ subbundle of $E$. The bundles $E^{\prime}$ and $E^{\prime\prime}$ are not necessarily holomorphic and hence they give a $C^{\infty}$ orthogonal decomposition $E=E^{\prime}\oplus E^{\prime\prime}$. However, there exists a proposition in complex geometry that guarantees that the above decomposition is also holomorphic if certain invariant condition holds. To be precise one has the following result.
Proposition 1.
([12], Prop. (4.18), p. 13). Let $(E,h)$ be a holomorphic hermitian vector bundle and $D_{h}$ the Chern connection of it. Let $E^{\prime}\subset E$ be a $D_{h}$-invariant $C^{\infty}$ complex subbundle and $E^{\prime\prime}$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Then both $E^{\prime}$ and $E^{\prime\prime}$ are $D_{h}$-invariant holomorphic subbundles of $E$ and they give a holomorphic orthogonal decomposition:
(1.1) $E=E^{\prime}\oplus E^{\prime\prime}.$
As it is well known, if $S\subset E$ is a rank $p$ holomorphic subbundle we have the following short exact sequence
(1.2) $0\longrightarrow S\longrightarrow E\longrightarrow Q\longrightarrow 0$
where $Q=E/S$ is a rank $r-p$ holomorphic vector bundle. Let $h$ be an hermitian metric in $E$ and $S^{\bot}$ the orthogonal complement of $S$ with respect to $h$, then $S^{\bot}$ is a $C^{\infty}$ complex vector bundle (not necessarily a holomorphic one) and hence $E=S\oplus S^{\bot}$ is –a priori– a $C^{\infty}$ complex orthogonal decomposition. As $C^{\infty}$ complex vector bundles $Q\cong S^{\bot}$ and the restriction of $h$ to $S^{\bot}$ induces –via this isomorphism– an hermitian metric $h_{Q}$ in $Q$. Therefore, if $h_{S}$ denotes the restriction of $h$ to $S$, the pairs $(S,h_{S})$ and $(Q,h_{Q})$ are both holomorphic hermitian vector bundles.
Given any $\xi\in A^{0}(S)$ we have a decomposition
(1.3) $D_{h}\xi=D_{h,S}\xi+A_{h}\xi\,,$
with $D_{h,S}\xi\in A^{1}(S)$ and $A_{h}\xi\in A^{1}(S^{\bot})$, and one has the following result.
Proposition 2.
([12], Prop. (6.4), p. 20). Let $(E,h)$ be a holomorphic hermitian vector bundle, where $D_{h}$ is the Chern connection of it and let $S\subset E$ be a holomorphic subbundle. Then, in the decomposition (1.3), the operator $D_{h,S}$ is the Chern connection of $(S,h_{S})$ and $A_{h}$ is a $(1,0)$-form with values in ${\rm Hom}(S,S^{\bot})$, i.e., $D_{h,S}=D_{h_{S}}$ and $A_{h}\in A^{1,0}({\rm Hom}(S,S^{\bot}))$.
In a similar form, for any $\eta\in A^{0}(S^{\bot})$ we have
(1.4) $D_{h}\eta=B_{h}\eta+D_{h,S^{\bot}}\eta\,,$
with $D_{h,S^{\bot}}\eta\in A^{1}(S^{\bot})$ and $B_{h}\eta\in A^{1}(S)$. Via the aforementioned isomorphism we can consider
(1.5) $D_{h,S^{\bot}}\equiv D_{h,Q}:A^{0}(Q)\longrightarrow A^{1}(Q)$
and one has the following result.
Proposition 3.
([12], Prop. (6.6), p. 21). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2 and $Q$ the holomorphic bundle given by (1.2). Then the operator $D_{h,Q}$ defined in (1.5) is the Chern connection of $(Q,h_{Q})$ and $B_{h}$ is a $(0,1)$-form with values in ${\rm Hom}(S^{\bot},S)$, i.e., $D_{h,Q}=D_{h_{Q}}$ and $B_{h}\in A^{0,1}({\rm Hom}(S^{\bot},S))$. Moreover, for any $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\bot})$ we have
(1.6) $h(A_{h}\xi,\eta)+h(\xi,B_{h}\eta)=0\,.$
The form $A_{h}$ in Proposition 2 is called the second fundamental form of $S$ in $(E,h)$ and again –via the isomorphism– one has $A_{h}\in A^{1,0}({\rm Hom}(S,Q))$ and $B_{h}\in A^{0,1}({\rm Hom}(Q,S))$. The expression (1.6) says that $A_{h}$ is the adjoint of $-B_{h}$ with respect to $h$. Now, as a consequence of the Leibniz rule, (1.3) and (1.4) can be naturally extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. [^1] Notice that if $\lambda\in A^{s}$, then $D_{h}(\lambda\xi)=(d\lambda)\xi+D_{h}\xi\wedge\lambda=D_{h,S}(\lambda\xi)+A_{h}\wedge(\lambda\xi)\,$. Therefore, for any $\xi\in A^{0}(S)$ we have
$R_{h}\xi$ $=$ $D_{h}(D_{h,S}\xi+A_{h}\xi)$
$=$ $D_{h,S}(D_{h,S}\xi)+A_{h}\wedge(D_{h,S}\xi)+B_{h}\wedge(A_{h}\xi)+D_{h,S^{\bot}}(A_{h}\xi)$
$=$ $(R_{h,S}+B_{h}\wedge A_{h}+D_{h}A_{h})\xi\,.$
Similarly, for any $\eta\in A^{0}(S^{\bot})$ we get
$R_{h}\eta=(R_{h,S^{\bot}}+A_{h}\wedge B_{h}+D_{h}B_{h})\eta\,.$
Putting all these together we can rewrite $R_{h}$ in a matrix form as follows:
(1.7) $R_{h}=\begin{pmatrix}R_{h,S}+B_{h}\wedge A_{h}&D_{h}B_{h}\\ D_{h}A_{h}&R_{h,S^{\bot}}+A_{h}\wedge B_{h}\end{pmatrix},$
which are Gauss-Codazzi type equations for holomorphic hermitian vector bundles (cf. [12], p. 23 or also [7], p. 274). Now, from Propositions 2 and 3 the forms $A_{h}$ and $B_{h}$ are of type $(1,0)$ and $(0,1)$ and since $R_{h}\in A^{1,1}({\rm End}E)$, the last terms in the right hand sides of $R_{h}\xi$ and $R_{h}\eta$ can be replaced by $d^{\prime\prime}A_{h}$ and $D^{\prime}_{h}B_{h}$. Consequently, the matrix expression (1.7) can be further simplified. Finally, one has the following result.
Proposition 4.
([12], Prop. (6.14), p. 23). Let $(E,h)$, $D_{h}$ and $S$ as in Proposition 2. If the second fundamental form $A_{h}$ of $S$ in $(E,h)$ vanishes identically, then the orthogonal complement $S^{\bot}\subset E$ is a holomorphic subbundle and
(1.8) $E=S\oplus S^{\bot}$
is a holomorphic orthogonal decomposition.
Proposition 4 is a straightforward consequence of Proposition 1. In fact, if $A_{h}$ vanishes identically, then (1.3) implies that $S\subset E$ is $D_{h}$-invariant. Notice that the converse of Proposition 4 is also true. In fact, if the decomposition $E=S\oplus S^{\bot}$ is holomorphic, then the pairs $(S,h_{S})$ and $(S^{\bot},h_{S^{\bot}})$ are both holomorphic hermitian vector bundles. If $D_{h_{S}}$ and $D_{h_{S^{\bot}}}$ are the Chern connections of these pairs, then $D_{h}$ is the direct sum of $D_{h_{S}}$ and $D_{h_{S^{\bot}}}$ and the result follows from (1.3) and Proposition 2, or equivalently from (1.4) and Proposition 3.
We would like to emphasize that propositions 2 and 3 have been already extended to Higgs bundles and appear –although using a local component notation– in the form of Proposition 2.3 in [2]. However, as far as the authors know, extensions of propositions 1 and 4 have not been explored yet in literature. As we already mentioned, it is the main purpose of this article to study in detail extensions of all these propositions to Higgs bundles.
2. Hermitian Higgs bundles and the second fundamental form
The aim of this section, is to revisit a couple of results on hermitian Higgs bundles previously known in literature [2] and which are extensions of propositions 2 and 3. We include here alternative proofs of these results without using local computations. In this section we also recall relevant information about Higgs bundles and fix the remaining part of the terminology that will be used throughout the article. In order to do this, we begin here reviewing some standard definitions on Higgs bundles (for more details the reader can see [5, 6] or the pioneering work [15]).
An hermitian Higgs bundle is a triple $({\mathfrak{E}},h)=(E,\Phi,h)$, where ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle [15] and $(E,h)$ is a holomorphic hermitian vector bundle [12]. In particular, the Higgs field $\Phi\in A^{1,0}({\rm End}E)$ is holomorphic and satisfies $\Phi\wedge\Phi=0$. Using the hermitian metric $h$ one has an hermitian conjugate ${\bar{\Phi}}_{h}\in A^{0,1}({\rm End}E)$ of the Higgs field and we get in the Higgs bundles setting the following connection:
(2.1) ${\cal D}_{h}=D_{h}+\Phi+{\bar{\Phi}}_{h}\,,$
usually known as the Hitchin-Simpson connection of $({\mathfrak{E}},h)$ [2]. A straightforward computation shows that its curvature ${\cal R}_{h}={\cal D}_{h}\wedge{\cal D}_{h}$ is given by
(2.2) ${\cal R}_{h}=R_{h}+D^{\prime}_{h}\phi+d^{\prime\prime}{\bar{\Phi}}_{h}+[\Phi,{\bar{\Phi}}_{h}]\,.$
Here, as we mentioned before, $R_{h}$ is the Chern curvature, $D^{\prime}_{h}$ and $d^{\prime\prime}$ are the $(1,0)$ and $(0,1)$ parts of $D_{h}$ and the last term is the usual graded commutator in the space of ${\rm End}E$-valued forms [6], i.e.,
$[\Phi,{\bar{\Phi}}_{h}]=\Phi\wedge{\bar{\Phi}}_{h}+{\bar{\Phi}}_{h}\wedge\Phi\,.$
If ${\mathfrak{E}}=(E,\Phi)$ and ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi^{\prime})$ are Higgs bundles (over the same $M$), a morphism of Higgs bundles $f:{\mathfrak{E}}\longrightarrow{\mathfrak{E}^{\prime}}$ is just a mapping $f:E\longrightarrow E^{\prime}$ such that
$\Phi^{\prime}\circ f=(f\otimes I)\circ\Phi\,,$
where $I$ denotes here the identity morphism of $\Omega^{1,0}$, i.e., it is a morphism of the corresponding holomorphic vector bundles in which the obvious square diagram commutes. If ${\mathfrak{E}}=(E,\Phi)$ is a Higgs bundle, a Higgs subbundle ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a $\Phi$-invariant holomorphic subbundle $S\subset E$. Hence, if $\Phi_{S}$ denotes the restriction of the Higgs field $\Phi$ to $S$, the pair $(S,\Phi_{S})$ is itself a Higgs bundle and the inclusion induces a morphism of Higgs bundles ${\mathfrak{S}}\longrightarrow{\mathfrak{E}}\,$.
The following is a well known result on Higgs bundles, however it plays a fundamental role in this article and hence we include it explicitly.
Lemma 1.
If ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, there exists a short exact sequence of Higgs bundles
(2.3) $0\longrightarrow{\mathfrak{S}}\longrightarrow{\mathfrak{E}}\longrightarrow{\mathfrak{Q}}\longrightarrow 0\,$
where ${\mathfrak{Q}}$ is the quotient Higgs bundle.
Proof.
Associated to $S\subset E$ we have the short exact sequence (1.2) with $Q=E/S$ holomorphic. By tensoring this sequence with $\Omega^{1,0}$ we get a commutative diagram with an induced Higgs field $\Phi_{Q}\in A^{1,0}(\mathrm{End}Q)$ for $Q$. Hence the pair $\mathfrak{Q}=(Q,\Phi_{Q})$ is a Higgs bundle and we get (2.3). ∎
Let $({\mathfrak{E}},h)$ be an hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ a Higgs subbundle. Since $S$ is $\Phi$-invariant it follows that $S^{\bot}$ is ${\bar{\Phi}}_{h}$-invariant. [^2] Let $\xi\in A^{0}(S)$ and $\eta\in A^{0}(S^{\bot})$. From the definition of hermitian conjugate and the $\Phi$-invariance of $S$ we get $h(\bar{\Phi}_{h}\eta,\xi)=h(\eta,\Phi\xi)=0$ and $\bar{\Phi}_{h}\eta\in A^{1}(S^{\bot})\,$. Moreover, from Section 1 we know that $Q\cong S^{\bot}$ (as $C^{\infty}$ complex vector bundles) and hence using Lemma 1 we get that $S^{\bot}$ is $\Phi$-invariant, which in turn implies that $S$ is $\bar{\Phi}_{h}$-invariant. From the above we immediately get the following result.
Lemma 2.
If ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, then the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant if and only if it is ${\cal D}_{h}$-invariant.
For future reference, we introduce the following terminology. If ${\mathfrak{S}}$ is a Higgs subbundle of ${\mathfrak{E}}$ and the holomorphic subbundle $S\subset E$ is $D_{h}$-invariant, we say that the Higgs subbundle $\mathfrak{S}$ is $D_{h}$-invariant. A similar definition applies to the Hitchin-Simpson connection ${\cal D}_{h}$.
If now $\xi\in A^{0}(S)$ and we use (2.1) and (1.3) we get a decomposition
${\cal D}_{h}\xi=(D_{h,S}+\Phi+\bar{\Phi}_{h})\xi+A_{h}\xi\,,$
with $(D_{h,S}+\Phi+\bar{\Phi}_{h})\xi\in A^{1}(S)$ and $A_{h}\xi\in A^{1}(S^{\bot})$. Notice that here we are using not only that $S$ is $\Phi$-invariant but also $\bar{\Phi}_{h}$-invariant. By defining
(2.4) ${\cal D}_{h,S}=D_{h,S}+(\Phi+\bar{\Phi}_{h})\lvert_{S}\,$
we can rewrite the above decomposition as:
(2.5) ${\cal D}_{h}\xi={\cal D}_{h,S}\xi+A_{h}\xi\,.$
The expression (2.5) is the natural extension to hermitian Higgs bundles of the classical decomposition for holomorphic vector bundles (1.3). From (2.4), (2.5) and Proposition 2 we immediately have the following result.
Proposition 5.
Let $({\mathfrak{E}},h)$ be an hermitian Higgs bundle, where ${\cal D}_{h}$ is the Hitchin-Simpson connection of it and let ${\mathfrak{S}}\subset{\mathfrak{E}}$ be a Higgs subbundle. Then the operator ${\cal D}_{h,S}$ defined in (2.4) is the Hitchin-Simpson connection of $({\mathfrak{S}},h_{S})$ and $A_{h}$ is the second fundamental form of Proposition 2, i.e., ${\cal D}_{h,S}={\cal D}_{h_{S}}$ and $A_{h}\in A^{1,0}({\rm Hom}(S,S^{\bot}))$.
This result can be seen as an extension to Higgs bundles of the classical Proposition 2. We would like to emphasize that, strictly speaking, Proposition 5 is already known in literature and it appears in the form of Proposition 2.3 in [2] (parts (i) and (ii)). However, in that reference the result is found and proved in a different way, in particular in that reference the authors introduce the main objects using local unitary frame fields. Here we only make explicit that Proposition 5 is closely related to a well known result in holomorphic hermitian vector bundles and that it can be also obtained without using a local computation.
On the other hand, if $\eta\in A^{0}(S^{\bot})$ and we use (2.1) and (1.4) we get a decomposition
${\cal D}_{h}\eta=B_{h}\eta+(D_{h,S^{\bot}}+\Phi+\bar{\Phi}_{h})\eta\,,$
with $(D_{h,S}+\Phi+\bar{\Phi}_{h})\eta\in A^{1}(S^{\bot})$ and $B_{h}\eta\in A^{1}(S)$. Notice that this time we are using that $S^{\bot}$ is $\Phi$-invariant and also $\bar{\Phi}_{h}$-invariant. By defining
(2.6) ${\cal D}_{h,S^{\bot}}=D_{h,S^{\bot}}+(\Phi+\bar{\Phi}_{h})\lvert_{S^{\bot}}\,$
we can rewrite the above decomposition as:
(2.7) ${\cal D}_{h}\eta=B_{h}\eta+{\cal D}_{h,S^{\bot}}\eta\,.$
The expression (2.7) is the natural extension to hermitian Higgs bundles of the classical decomposition for holomorphic hermitian vector bundles (1.4). Now, via the aforementioned isomorphism and using (1.5) we can consider
(2.8) ${\cal D}_{h,S^{\bot}}\equiv{\cal D}_{h,Q}:A^{0}(Q)\longrightarrow A^{1}(Q)$
and from (2.6)-(2.8), Proposition 3 and Lemma 1 we immediately get the following result.
Proposition 6.
Let $({\mathfrak{E}},h)$, ${\cal D}_{h}$ and ${\mathfrak{S}}$ as in Proposition 5 and ${\mathfrak{Q}}$ the quotient Higgs bundle given by Lemma 1. Then the operator ${\cal D}_{h,Q}$ defined in (2.8) is the Hitchin-Simpson connection of $({\mathfrak{Q}},h_{Q})$ and $B_{h}$ is the same form of Proposition 3, i.e., ${\cal D}_{h,Q}={\cal D}_{h_{Q}}$ and $B_{h}\in A^{0,1}({\rm Hom}(S^{\bot},S))$.
The above proposition can be seen as a Higgs extension of the classical Proposition 3. Notice that the form $B_{h}$ is mentioned as part (iii) of Proposition 2.3 in [2]. The remaining content of this result is already implicit in the local computations in that reference. Again, here we only make explicit that Proposition 6 is closely related to a well known result in holomorphic hermitian vector bundles and that it can be obtained without using a local computation.
We would like to emphasize that the forms $A_{h}$ and $B_{h}$ of propositions 5 and 6 are the same as the classical forms for holomorphic subbundles of hermitian holomorphic bundles, a fact also mentioned in [2]. Therefore, if $({\mathfrak{E}},h)$ is an hermitian Higgs bundle and ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle, the second fundamental form of ${\mathfrak{S}}$ in $({\mathfrak{E}},h)$ is just the second fundamental form of the holomorphic subbundle $S$ in $(E,h)$. Now, as in the case of holomorphic hermitian vector bundles, (2.5) and (2.7) can be extended to decompositions for $s$-forms evaluated on $S$ and $S^{\bot}$, respectively. Therefore, for any $\xi\in A^{0}(E)$ we get
${\cal R}_{h}\xi=({\cal R}_{h,S}+B_{h}\wedge A_{h}+{\cal D}_{h}A_{h})\xi\,,$
and if $\eta\in A^{0}(S^{\bot})$ we obtain
${\cal R}_{h}\eta=({\cal R}_{h,S^{\bot}}+A_{h}\wedge B_{h}+{\cal D}_{h}B_{h})\eta\,.$
As it is usual, the above expressions can be used to rewrite ${\cal R}_{h}$ in a matrix form as:
(2.9) $\mathcal{R}_{h}=\begin{pmatrix}\mathcal{R}_{h,S}+B_{h}\wedge A_{h}&\mathcal{D}_{h}B_{h}\\ \mathcal{D}_{h}A_{h}&\mathcal{R}_{h,S^{\bot}}+A_{h}\wedge B_{h}\end{pmatrix},$
which is the Higgs extension of the Chern curvature matrix formula (1.7). The expression (2.9) can be considered as the Gauss-Codazzi type equations for hermitian Higgs bundles (cf. [2], p. 7).
3. Orthogonal decompositions of hermitian Higgs bundles
The aim of this section is to establish extensions of propositions 1 and 4 to the Higgs bundles setting. In particular, we show that it is possible to prove such extended results on Higgs bundles using definitions and results on (holomorphic) Higgs bundles. Hence, strictly speaking, the above extensions only depend on the standard notions given in the previous sections.
According to the terminology introduced in sections 1 and 2 we are now in position to establish the following result.
Proposition 7.
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle and $D_{h}$, $E^{\prime}$ and $E^{\prime\prime}$ as in Proposition 1 and suppose that $E^{\prime}$ is $\Phi$-invariant. Then both ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ and ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ are $D_{h}$-invariant Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and they give a Higgs orthogonal decomposition:
(3.1) ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}.$
Proof.
From Proposition 1 we know that $E^{\prime},E^{\prime\prime}\subset E$ are $D_{h}$-invariant holomorphic subbundles and we have the holomorphic orthogonal decomposition (1.1), i.e., $E=E^{\prime}\oplus E^{\prime\prime}$ in the holomorphic category.
Since $E^{\prime}$ is holomorphic and also $\Phi$-invariant, the pair ${\mathfrak{E}}^{\prime}=(E^{\prime},\Phi\lvert_{E^{\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, by defining ${\mathfrak{S}}={\mathfrak{E}}^{\prime}$ it follows that ${\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle and hence $S^{\bot}=E^{\prime\prime}$ is also $\Phi$-invariant. Since $E^{\prime\prime}$ is holomorphic, it follows that ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$ is a $D_{h}$-invariant Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (3.1). ∎
The above result can be seen as an extension to Higgs bundles of Proposition 1. More precisely, Proposition 7 guarantees that the same classical hypothesis, together with an additional $\Phi$-invariance condition, implies that ${\mathfrak{E}}^{\prime},{\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ are Higgs subbundles and that (3.1) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles.
Proposition 8.
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle, $D_{h}$ the Chern connection of $(E,h)$ and $S\subset E$ a holomorphic subbundle. If the second fundamental form $A_{h}$ of $S$ on $(E,h)$ vanishes identically and $S$ is $\Phi$-invariant, then ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ and ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ are Higgs subbundles of ${\mathfrak{E}}=(E,\Phi)$ and
(3.2) ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$
is a Higgs orthogonal decomposition.
Proof.
From Proposition 4 we know that $S^{\bot}\subset E$ is a holomorphic subbundle and we have the holomorphic orthogonal decomposition (1.8), i.e., $E=S\oplus S^{\bot}$ in the holomorphic category.
Since $S\subset E$ is holomorphic and $\Phi$-invariant, the pair ${\mathfrak{S}}=(S,\Phi\lvert_{S})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$. In particular, $S^{\bot}$ is $\Phi$-invariant and since it is also holomorphic, the pair ${\mathfrak{S}}^{\bot}=(S^{\bot},\Phi\lvert_{S^{\bot}})$ is a Higgs subbundle of ${\mathfrak{E}}=(E,\Phi)$ and we obtain (3.2). ∎
The above proposition can be seen as a Higgs extension of Proposition 4. In particular, Proposition 8 guarantees that the same classical hypothesis, together with an additional $\Phi$-invariance condition, implies that ${\mathfrak{S}},{\mathfrak{S}}^{\bot}\subset{\mathfrak{E}}$ are Higgs subbundles and that (3.2) is a decomposition of the original Higgs bundle in terms of these Higgs subbundles. It is important to mention that this proposition can also be proved using Proposition 7 and Lemma 2. In fact, by redefining $E^{\prime}=S$ and $E^{\prime\prime}=S^{\bot}$ and since $E^{\prime}$ is $\Phi$-invariant, then ${\mathfrak{E}}^{\prime}={\mathfrak{S}}\subset{\mathfrak{E}}$ is a Higgs subbundle. If $A_{h}$ vanishes identically, (2.5) implies that $E^{\prime}$ is ${\cal D}_{h}$-invariant. Therefore, from Lemma 2 we know that $E^{\prime}$ is also $D_{h}$-invariant and the result follows from Proposition 7.
As in the classical case, the converse of Proposition 8 is also true. In fact, if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs decomposition, the pairs $({\mathfrak{S}},h_{S})$ and $({\mathfrak{S}}^{\bot},h_{S^{\bot}})$ are both hermitian Higgs bundles. If ${\cal D}_{h_{S}}$ and ${\cal D}_{h_{S^{\bot}}}$ are the Hitchin-Simpson connections of these pairs, then ${\cal D}_{h}$ is the direct sum of ${\cal D}_{h_{S}}$ and ${\cal D}_{h_{S^{\bot}}}$ and the result follows from (2.5) and Proposition 5, or equivalently from (2.7) and Proposition 6.
We are now in position to mention some applications of the above results in the Higgs bundles setting. Recall that in Section 2 we have mentioned that the second fundamental form of a Higgs subbundle of a hermitian Higgs bundle is the same as the second fundamental form of the corresponding holomorphic subbundle. From this fact, the previous paragraph and Proposition 8, we immediately get the following result.
Corollary 1.
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle, ${\cal D}_{h}$ the Hitchin-Simpson connection of $({\mathfrak{E}},h)$ and ${\mathfrak{S}}\subset{\mathfrak{E}}$ a Higgs subbundle. Then, the second fundamental form $A_{h}$ of ${\mathfrak{S}}$ in $({\mathfrak{E}},h)$ given by (2.5) vanishes identically if and only if ${\mathfrak{E}}={\mathfrak{S}}\oplus{\mathfrak{S}}^{\bot}$ is a Higgs orthogonal decomposition.
We would like to emphasize that this result can be completely established in terms of objects in the Higgs category, therefore it is natural to wonder if there exists another proposition, closely related to Proposition 7, and such that can be established making reference only to objects in the Higgs category. As we will see in a moment, it is indeed the case and Lemma 2 is a key result in the proof of such a proposition. Before doing that, let us introduce the following definition. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a Higgs subbundle and $E^{\prime\prime}\subset E$ the orthogonal complement of $E^{\prime}$ with respect to $h$. Let us consider the pair ${\mathfrak{E}}^{\prime\prime}=(E^{\prime\prime},\Phi\lvert_{E^{\prime\prime}})$, which we call the $C^{\infty}$ Higgs orthogonal complement of the Higgs subbundle ${\mathfrak{E}}^{\prime}$ with respect to $h$. The bundle $E^{\prime\prime}\subset E$ is just a $C^{\infty}$ complex subbundle (not necessarily holomorphic) and hence, strictly speaking, ${\mathfrak{E}}^{\prime\prime}$ is not necessarily a Higgs subbundle of ${\mathfrak{E}}$. However, that will be the case if certain invariant condition holds. To be precise, we have the following result.
Corollary 2.
Let $(\mathfrak{E},h)=(E,\Phi,h)$ be an hermitian Higgs bundle and ${\cal D}_{h}$ the Hitchin-Simpson connection of $({\mathfrak{E}},h)$. Let ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ be a ${\cal D}_{h}$-invariant Higgs subbundle and ${\mathfrak{E}}^{\prime\prime}$ the $C^{\infty}$ Higgs orthogonal complement of ${\mathfrak{E}}^{\prime}$ with respect to $h$. Then ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ is also a ${\cal D}_{h}$-invariant Higgs subbundle and ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ is a Higgs orthogonal decomposition.
Proof.
By applying Lemma 2 to the Higgs subbundle ${\mathfrak{E}}^{\prime}\subset{\mathfrak{E}}$ we know that $E^{\prime}$ is $D_{h}$-invariant. Since it is also $\Phi$-invariant, Proposition 7 implies the Higgs orthogonal decomposition ${\mathfrak{E}}={\mathfrak{E}}^{\prime}\oplus{\mathfrak{E}}^{\prime\prime}$ with ${\mathfrak{E}}^{\prime\prime}\subset{\mathfrak{E}}$ also a $D_{h}$-invariant Higgs subbundle. The ${\cal D}_{h}$-invariance of this last subbundle follows again from Lemma 2. ∎